PLEASE SHOW ALL WORK THANKSMicrophone - - - - Speakers Click to go forward, hold to see history Two small speakers are separated by a distance of 6 cm,
Answer #1

Frequency = 14kHz = 14000 Hz

wavelength of the signal will then be: \lambda = \frac{c}{f} = 0.0245 m

For 1st order minima,

d\frac{x}{L} = \frac{\lambda}{2}

0.06\frac{x}{1.4} = \frac{0.0245}{2}

=> x = 0.286m

this is the distance from the center where the first minima occurs.